OpenStudy (anonymous):

The sum of the reciprocals of two consecutive integers is 19/90. Find the two integers.

OpenStudy (anonymous):

Two consecutive integers are x and x+1 Their reciprocals are 1/x and 1/(x+1) 1/x + 1/(x+1) = 19/90 Multiply both sides by 90x(x+1) 90(x+1) + 90x = 19x(x+1) 90x + 90 + 90x = 19x² + 19x 19x² + 19x - 180x - 90 = 0 19x² - 161x - 90 = 0 19x² - 171x + 10x - 90 = 0 19x (x - 9) + 10 (x - 9) = 0 (x - 9) (19x + 10) = 0 x = 9, x = -10/19 But -10/19 is not an integer So the two integers are 9 and 10

OpenStudy (anonymous):

Medal?